dejdaddy45:

Find the inverse of f(x) for f(x) = -3x +2 and state whether or not it is a function. a. inv f(x) = -(1/3)x + 2/3; function c. inv f(x) = (1/3)x + 2/3; function b. inv f(x) = (1/3)x - 2/3; not a function d. inv f(x) = -(1/3)x - 2/3; function

Narad:

Let \[y=-3x+2\] Then, \[3x=-y+2\] \[x=-1/3y+2/3\] Therefore, \[f ^{-1}(x)=-1/3x+2/3\] The answer is option (a). This is a function